OpenStudy (anonymous):

Solve: x2 – 2x = 15 A. {-3, 5} B. {-5, -3} C. {3, 5} D. {-5, 3}

OpenStudy (king):

u can split the middle term do u know hw to split the middle term....

mathslover (mathslover):

\[\huge{x^2-2x-15}\] \[\huge{x^2-5x+3x-15}\] can u factorize this ?

OpenStudy (king):

yeah maths lovers rite thatz hw u split the middle term do u know hw to do ahead? @senior012

mathslover (mathslover):

@Senior012 you must focus on how to take terms common i mean how to factor an expression: \[\huge{ab+ac=a(b+c)}\] or what about this example ? \[\huge{a^2-ab=a(a-b)}\]

mathslover (mathslover):

@Senior012 r u getting that all . or shall i try to explain u in an easier way ?

OpenStudy (anonymous):

A is correct answer.